OpenStudy (anonymous):

can someone help me evaluate the limit of this equation limit approaches 0(e^(4x) - 1-4x)/(x^2)

OpenStudy (anonymous):

0

OpenStudy (anonymous):

i get zero over zero

OpenStudy (amistre64):

derive top and bottom seperately; then limit them seperately

OpenStudy (amistre64):

lhopitals rule

OpenStudy (amistre64):

i wonder if you can do that more than once

OpenStudy (anonymous):

Yes. You can do it as long as both the denominator and numerator have 0/0 or ∞/∞

OpenStudy (amistre64):

2x doesnt seem to help does it lol

OpenStudy (amistre64):

you can? over and over?

OpenStudy (anonymous):

yeah as long as you get 0/0 or infin/infin

OpenStudy (amistre64):

i thought i read that somewhere lol

OpenStudy (anonymous):

ok so i would get the derivative of the top correct over the derivative of the bottom

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

4.e^4x - 4 2.e^4x - 2 ---------- = --------- 2x x

OpenStudy (amistre64):

8.e^4x at 0 = ...8 right?

OpenStudy (anonymous):

yeah thats what the answer sheet got but i still dont understand how you got it

OpenStudy (amistre64):

x^2; 2x ; 2 tese are the derivatives right? of the bottom

OpenStudy (anonymous):

yeah

OpenStudy (amistre64):

e^4x -1 - 4x; 4.e^4x -4; 16.e^4x for the top

OpenStudy (amistre64):

16.e^4x ------- = 8.e^4x ; at x=0; 8.1 = 8 2

OpenStudy (anonymous):

cant see where you get the 16 from sorry

OpenStudy (amistre64):

4.e^4x right? derive e^4x again and we get: 4. 4.e^4x = 16.e^4x

OpenStudy (amistre64):

e^u derives to : Du e^u

OpenStudy (anonymous):

ok so u used l hopital twice

OpenStudy (amistre64):

yes lol

OpenStudy (anonymous):

ok ok awesome and we were just talking about that